Had problems initiallizing the COSMOS communications with the IEB during the afternoon. We followed up on Jay Elias' suggestion to check the ethernet cables. The ethernet cable from the IEB was plugged into port 11 on the switch in the cass cage. COSMOS will only work if this ethernet cable is plugged into port 17.
Point was bad at the beginning of the night. The points was fixed so guiding tests were begun. Afterwards a we took data for a new pointing map. Rolo, Roberto, and Omar will have more details about this.
Once the pointing map was made, we were ready to begin with COSMOS. We noticed that when using the comparison lamp system that the detector is not uniformly illuminated by the HgAr penray or the CuHeAr hollow cathode. Nevertheless, we were able to take camera focus sequences for the spectroscopic setups. We then took some test dome flats to check the illumination of the detector. This worked well and the spectroscopic flats look normal.
Contrary to previous discussions I had with Javier, Manuel H., and Manuel M., I will not be using the ADC tomorrow night.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-05-05 | 0.0 | 11.25 | 0.0 | 0.0 | 0.0 | 11.25 |
None.